10. How much energy must be added to 2800 g of gold at its melting point...

10.
How much energy must be added to 2800 g of gold at its melting point of 1063 deg. C in order to change it from a solid to a liquid at 1121.0 deg. C?

Answer:  Last Answer: 1.7676E+05 J
Not yet correct, tries 3/20

I NEED THE UNITS TO BE CORRECT PLEASE. THEY'RE BEING USED ON CAPA. ALSO SHOW ALL WORK. THANKS.

Thermal Properties of Some Materials

Material alpha beta csolid Tmelt Lfusion cliquid Tboil Lvaporize cgas k
K-1 K-1 J kg-1 K-1 oC J kg-1 J kg-1 K-1 oC J kg-1 J kg-1 K-1 J s-1 m-1 K-1
Air -- -- -- -- -- -- -- -- -- 0.0256
Aluminum 26 x 10-6 75 x 10-6 900 660 380 x 103 900 2450 11400 x 103 900 240
Ammonia -- -- 4700 -77.8 3320 x 103 4700 -33.4 1371 x 103 2060 --
Brass 19 x 10-6 C-1 56 x 10-6 C-1 -- -- -- -- -- -- -- 110
Concrete -- -- -- -- -- -- -- -- -- 1.1
Copper 17 x 10-6 C-1 51 x 10-6 C-1 378 1083 134 x 103 378 2565 5069 x 103 378 390
Ethyl Alcohol -- 1100 x 10-6 C-1 2450 -114.4 1080 x 103 2450 78.3 855 x 103 1680 --
Gasoline 0 x 10-6 C-1 950 x 10-6 C-1 -- -- -- -- -- -- -- --
Glass 9 x 10-6 C-1 27 x 10-6 C-1 -- -- -- -- -- -- -- 0.8
Ethelyne Glycol 0 x 10-6 C-1 570 x 10-6 C-1 -- -- -- -- -- -- -- --
Gold 14 x 10-6 C-1 42 x 10-6 C-1 129 1063 64.5 x 103 129 2660 1578 x 103 129 --
Goose Down -- -- -- -- -- -- -- -- -- 0.025
Hydrogen -- -- 967 -259.3 58.6 x 103 967 -252.9 452 x 103 1432 --
Invar 0.9 x 10-6 C-1 2.7 x 10-6 C-1 -- -- -- -- -- -- -- --
Iron 12 x 10-6 C-1 35 x 10-6 C-1 -- -- -- -- -- -- -- 79
Kerosene 0 x 10-6 C-1 990 x 10-6 C-1 -- -- -- -- -- -- -- --
Lead 29 x 10-6 C-1 87 x 10-6 C-1 128 327 24.5 x 103 128 1750 871 x 103 128 35
Nitrogen -- -- 2042 -210 25.5 x 103 2042 -195.8 201 x 103 1040 --
Oxygen -- -- 1669 -218.8 13.8 x 103 1669 -183 213 x 103 919 --
Pyrex 3 x 10-6 C-1 9 x 10-6 C-1 -- -- -- -- -- -- -- --
Silver 18 x 10-6 C-1 18 x 10-6 C-1 -- -- -- -- -- -- -- 420
Steel 12 x 10-6 C-1 35 x 10-6 C-1 449 1538 -- -- 2862 -- -- 79
Styrofoam -- -- -- -- -- -- -- -- -- 0.01
Water -- 214 x 10-6 C-1 211 0 334 x 103 4186 100 2256 x 103 2080 0.6
Wood -- -- -- -- -- -- -- -- -- 0.15
Wool -- -- -- -- -- -- -- -- -- 0.04
Answer #1

Mass of gold M=2800g

Initial temperature T1=1063 °C

finial temperature T2=1121°C

Specific heat s=129

This concept belongs thermal properties of meter

So heat energy required for specific heat

Q=MS(t2-t1)

=2800×129(1121-1063)

=20949600J

=209.5×10^5 J

Therefore the heat energy required =209.5×10^5 J